OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

y-y0=m(x-x0) ==> according this we have: y-(-3)=-4(x+8) ==> y+3)=-4(x+8)